When setting the price of admission tickets for the amusement park, Tabitha wants the price for a two-child, two adult family to be $100. She also wants the adult tickets to cost $8.00 more than twice the cost of a child's ticket. How much should a child's ticket cost?
Answer by rfer(16322)

2x+2(2x+8)=100
2x+4x+16=100
6x=84
x=$14
2x+8=$36